OpenStudy (anonymous):

Apples sell for $1.90 per pound, and bananas sell for $0.75 per pound. Troy bought some apples and some bananas. Together they weighed 3.8 pounds, and cost $5.84. How many pounds of apples and how many pounds of bananas did Troy buy? A. 1.5 pounds of apples; 2.3 pounds of bananas B. 2.6 pounds of apples; 1.2 pounds of bananas C. 1.2 pounds of apples; 2.6 pound of bananas D. 1.9 pounds of apples; 1.9 pounds of bananas

OpenStudy (anonymous):

\[x+y=3.8\] \[1.9x+.75y=5.84\]

OpenStudy (anonymous):

solve for an x or y and substitute

OpenStudy (anonymous):

i would go y= 3.8-x substitute that y into 1.9x+.75(3.8-x)=5.84

OpenStudy (anonymous):

that will give you x pounds to substitute into x+y=3.8

OpenStudy (anonymous):

1.15p + 2.85 = 5.84 is the same as which of these? 1.5 pounds of apples; 2.3 pounds of bananas 2.6 pounds of apples; 1.2 pounds of bananas 1.2 pounds of apples; 2.6 pound of bananas 1.9 pounds of apples; 1.9 pounds of bananas

OpenStudy (anonymous):

can you figure it out with the above responses?

OpenStudy (anonymous):

To difficult

OpenStudy (anonymous):

1.9x+.75(3.8-x)=5.84

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

can you solve that for x

OpenStudy (anonymous):

brb ill check your answer if you like